Information regarding tasks in an assembly line run for an 8-hour shift each day is provided:

Task

Proceeding Task (s)

Tme (min)

A

-

5

B

-

3

C

-

2

D

A

1

E

B

4

F

C

6

G

D,E

7

H

G

5

I

F,G

6

J

H,1

4

K

I

3

a)What is the theoretical minimum number of stations for a shift output of 40 units?

B)if the line is balanced according to the desired cycle time and with one more than the minimum number of stations wht would be % idle time?
Answer #1

Output = 40 units

Shift time = 8 hours or 480 minutes

Cycle time= Shift time / Output = 480/40 = 12 minutes

Sum of all task times = 46 minutes

Theoretical minimum number of stations = 46 / 12 = ~4 workstations

Idle time when there are 4+1 = 5 Workstations = Sum of task time / Number of workstations * Cycle time = 46 / (5*12) = 46/60 = 76.67%